# In-context Learning for Automated Driving Scenarios

### v1

@@ Abstract
One of the key challenges in current Reinforcement Learning (RL)-based Automated Driving (AD) agents is achieving flexible, precise, and human-like behavior cost-effectively. This paper introduces an innovative approach utilizing Large Language Models (LLMs) to intuitively and effectively optimize RL reward functions in a human-centric way. We developed a framework where instructions and dynamic environment descriptions are input into the LLM. The LLM then utilizes this information to assist in generating rewards, thereby steering the behavior of RL agents towards patterns that more closely resemble human driving. The experimental results demonstrate that this approach not only makes RL agents more anthropomorphic but also reaches better performance. Additionally, various strategies for reward-proxy and reward-shaping are investigated, revealing the significant impact of prompt design on shaping an AD vehicle’s behavior. These findings offer a promising direction for the development of more advanced and human-like automated driving systems. Our experimental data and source code can be found here [^1] GitHub: https://github.com/JingYue2000/In-context_Learning_for_Automated_Driving..
@@
Large language models, Automated driving, Reinforcement learning, DQN, In-context learning
@@ I Introduction
With the advancement of artificial intelligence, an increasing number of machine learning techniques are being applied in the field of intelligent transportation [1][2]. Among them, Reinforcement Learning (RL) is a subfield of machine learning that deals with how an agent can learn to make decisions and take actions in an environment to achieve specific goals following a trial-and-error learning approach [3]. Reinforcement Learning showcases robust learning and adaptability, particularly in real-time decision-making scenarios prevalent in automated driving [4, 5, 6].
@@
  • Perfecting reinforcement learning agents to mirror human driving while ensuring safety and efficiency remains a complex task.
@@
To solve these challenges, a shift in approach is emerging with In-context learning (ICL),  [10], where LLMs make predictions only based on contexts augmented with a few examples [11]. By specifying an objective with a natural language prompt, LLM can convert user goals into reward signals, and guide RL agents intuitively toward desired outcomes [12]. This innovative interaction between LLMs and RL agents not only advances opportunities in fields like automated driving but also highlights the LLMs’ versatility in complex decision-making.
Building on this innovation, we introduce our approach for leveraging the synergy between LLMs and RL agents in the context of automated driving scenarios. We explore the mechanisms by which the reward guidance offered by LLMs influences the training and performance of RL agents. Further, we explore methodologies to achieve diverse driving behaviors through the integrated LLM-RL framework, demonstrating the potential of this collaboration in enhancing autonomated driving technologies.
The contributions of this paper are summarized as follows:
  • We proposed an RL training framework incorporating LLMs to simplify RL-AD agents’ reward design in AD scenarios, quoted as LLM-RL agent in this paper.
@@
The remainder of this paper is organized as follows. Section II provides an overview of the related work followed by presenting the problem formulation and approach in III. The proposed approach is further described in Section IV. The experiments and results are discussed in Section V. Finally, the conclusion and future work are summarized in Section VI.
@@ II-A Large Language Models in Reinforcement Learning
Large Language Models (LLMs), known for their capacity to learn from minimal context and deliver human-like responses following training with human data [13], are increasingly considered as a potential solution when combined with Reinforcement Learning (RL) agents to perform higher-level tasks. In certain studies, RL is employed to fine-tune existing LLMs[14, 15, 16]. Conversely, other research integrates RL with LLMs to execute specific tasks or develop particular skills[17, 18, 19]. However, our research diverges from these methods by using LLMs as proxy rewards, enhancing RL model training as described in[20].
The research landscape in using large language models (LLMs) as proxies for reinforcement learning (RL) rewards has varied approaches. Kwon et al. pioneered the use of LLMs for learning in contexts such as the Ultimatum Game and negotiation tasks, focusing solely on LLM-generated rewards. [12]. In contrast, our study utilizes a diverse set of rewards in dynamic AD system scenarios, broadening the scope of possibilities. Other researchers, like Xie et al. and Ma et al. [21][22], have explored reward design through direct Python code generation. Meanwhile, Song et al. in [23] required human inputs to iteratively fine-tune rewards, whereas another study [24] used LLM outputs as auxiliary rewards. Unlike these approaches, our method directly employs LLM responses to generate reward signals, simplifying the process and eliminating the need for human intervention.
@@ II-B Large Language Models in Automated Driving
Utilizing Large Language Models (LLMs) in automated driving enhances decision-making by tapping into LLMs’ vast knowledge and reasoning capabilities. Research in vehical planning and control mainly falls in performing Prompt Engineering  [25, 26]or Fine-tuning Pre-trained Model as described in  [27, 28].
Targeting at solving closed-loop driving tasks, Fu et al.  [26] utilize LLMs to comprehend and interact with environments. Wen et al. improved this framework and help LLM make decision with different reflection modules. [25]. Our approach simplifies decision-making by considering only the ego car’s most recent action. Similarly, Cui et al.  [29] proposed a method that can provide the LLM with comprehensive environment information from data collected from different modules of the vehicles, ensuring safer results from LLM decisions. Despite these advancements, these approaches suffer from real-time latency in decision-making.
Other researchers focused on fine-tuning the pre-trained model, like Chen et al.  [28] performed fine-tuning with 10,000 generated driving scenarios. DriveMLM [30] merged linguistic processing with vehicle control for a more advanced system. Our work leverages LLM to enhance RL performance, satisfying real-time requirements by dynamically querying it based on scenarios.
@@ III-A Markov Decision Process Formalization
An automated driving scenario can be modeled as a Markov Decision Process: $M=\langle S,A,\rho,R,\gamma\rangle$, where $S$ is the state space representing various driving scenarios (e.g., traffic density, relative speeds of vehicles, lane positions, etc.). $A$ is the action space representing the set of all possible driving actions (e.g., accelerate, decelerate, change lanes, etc.). $\rho:S\times A\times S\rightarrow[0,1]$ is the transition probability function representing the likelihood of transitioning from one state to another given an action. $R:S\times A\rightarrow\mathbb{R}$ is the traditional reward function mapping states and actions to real-valued rewards. $\gamma$ is the discount factor, indicating the agent’s consideration for future rewards.
Inspired by [12] we employed a LLM as the proxy for reward calculations in our framework to achieve a more interpretable and adaptable reward mechanism.
@@
As illustrated in Fig. 1 our integrated framework uses the LLM to interpret textual prompts $\varrho$ as input and outputs a reward signal. The text prompt $\varrho$ consists of three major components: Task Description $\rho 1$, User Objective $\rho 2$ and Last Outcome $\rho 3$. The LLM is modeled as a function LLM: $A*\rightarrow A*$, taking the concatenated prompt $\varrho$ as input and outputting a string. A parser $A*\rightarrow{0,1}$ is then used to map the textual output of the LLM to a binary reward signal, which refers to the (2) of Fig. 1. Then, the binary reward signal will be sent to the RL agent for further training, which refers to the (3) of Fig. 1. This framework replaces the traditional reward function $R$ with a proxy reward function LLM and can be used with any RL training algorithm.
@@ III-B Reward Function Shaping Strategies
LLM’s response subsequently translated into a quantifiable reward signal. This reward, termed "LLM Reward" in this paper, is an integer, which is 0 or 1, reflecting the appropriateness of the agent’s actions as evaluated by the LLM. This innovative approach aims to leverage the nuanced understanding capabilities of LLMs to refine and guide the RL agent’s decision-making process.
We used combined reward to take advantage of both Deep Q Learning (DQN) and LLM, referring to the method mentioned in [31]. Given the original reward function in a highway driving environment, we propose an adjusted reward function that integrates a reward signal from LLM and scales the combined reward to the [0,1] range using a tangent function:
$R_{\text{ad}}(s,a)=\sigma\left(\text{Env\_reward}+\cdot\text{LLM\_reward}(s,a)\right)$ (1)
@@ IV Proposed Approach
A standard DQN model is set up as a baseline agent, trained with reward function 2 that are calculated based on the current driving scenario, which serves as a foundation of our proposed approach.
@@
These prompts distinctly outline the scenario as well as instruction for LLM to make decision for further steps.
Scenario informations included are: speed of the vehicle in control(the LLM-RL agent), speed and positions of nearby vehicles, LLM-RL agent’s last action and possible actions to choose from.
Further instructions including driving rules, caution points and specified format needed for generating and presenting responses.
@@ IV-A Basic Driving Rules and Task Description
To illustrate the context for the ChatGPT agent, we provide a series of predefined prompts:
  • SYSTEM_MESSAGE: Specifies LLM’s role and setting the scenario’s context.
  • Driving_RULES: Provides the general safe-driving guidelines which shapes the behaviour pattern for the driving styles of the agents.
@@
These predefined messages and rules can serve as a foundation for the behaviour of the driving assistant. The construction and phrasing of previous descriptions are based on the work of drive-like-human [33].
@@ IV-B Context Description
To illustrate the context of the current road situation, the following basic elements are defined as Lane and Vehicle. The Vehicle class encapsulates properties including speed in both directions and positional data, while the Lane class incorporates attributes including position, unique identification, and details regarding adjacent lanes.
@@ IV-D Formatting Output
This part of the prompt aims at formatting an output string that is interpretable to our program during runtime. We specify the outputs of the agent to generate only the following five actions of the ego vehicle: LANE_LEFT, IDLE, LANE_RIGHT, FASTER, and SLOWER. The LLM interprets all information above context and generates feedback. A procedure to generate a final reward is discussed as follows:
We initialize Highway_Env and input a concatenated prompt, which includes the last action, the pre-define prompt, and the scene description. The last action is derived from the previous decision made by the agent. This prompt enables the agent to understand the scenario it faces and ensures that the responses generated by the agent adhere to user requirements.
@@
V Experiments and Analysis
@@ V-A Highway AD as a case study
The practical capabilities of leveraging the in-context learning ability of LLM in AD are illustrated through highway driving scenarios. The potential for AD agent personalization in highway scenarios by establishing two groups with different meta-prompts is also explored in this case study. In the prompt case shown in Fig. 3, the scenario can be described as the ego car driving in lane 1 on a highway. There is one vehicle in its current lane. The right lane change is deemed safe by the safety assessment. The previous decision the ego car made, was to stay idle, which means the ego car is in the process of following the front car. This scenario featured two distinct driving agents: the "aggressive" and "conservative" agents. A comprehensive analysis was conducted under controlled conditions to compare their performance.
Figure 3: Meta prompt case study
Each RL agent’s driving behaviour was shaped through a series of user instructions, establishing their respective driving style within the scenario as shown in Fig. 4. Agent’s decisions show strong consistency in humanized behavior reflecting the behavior pattern meta-prompt.
The agent consider several factors: Safety, Efficiency and Driving Rules. For an aggressive agent, the user has stated that efficiency is of higher priority, hence the agent is considering lane right for higher efficiency. For a conservative agent, the user has stated that safety is of higher priority, hence the agent is considering slowing down to keep distance from the leading car.
@@
Compared with manually designing a reward function, this manner is simpler and has demonstrated its effectiveness in guiding the behaviour of an agent. These observations indicate that Large language models (LLMs), such as GPT in our case, offer a potential cost-efficiency alternative to performing similar roles with human feedback in certain contexts. LLMs can leverage vast amounts of text data to learn implicitly from human-generated content, interactions, and feedback, significantly reducing the human workload in reward design while achieving similar effects.

### v2

# Human-centric Reward Optimization for Reinforcement Learning-based Automated Driving using Large Language Models

@@ Abstract
One of the key challenges in current Reinforcement Learning (RL)-based Automated Driving (AD) agents is achieving flexible, precise, and human-like behavior cost-effectively. This paper introduces an innovative approach that uses large language models (LLMs) to intuitively and effectively optimize RL reward functions in a human-centric way. We developed a framework where instructions and dynamic environment descriptions are input into the LLM. The LLM then utilizes this information to assist in generating rewards, thereby steering the behavior of RL agents towards patterns that more closely resemble human driving. The experimental results demonstrate that this approach not only makes RL agents more anthropomorphic but also achieves better performance. Additionally, various strategies for reward-proxy and reward-shaping are investigated, revealing the significant impact of prompt design on shaping an AD vehicle’s behavior. These findings offer a promising direction for the development of more advanced, human-like automated driving systems. Our experimental data and source code can be found here [^1] GitHub: https://github.com/JingYue2000/In-context_Learning_for_Automated_Driving..
@@
Large language models, automated driving, reinforcement learning, DQN, PPO, reward function optimization, in-context learning.
@@ I Introduction
With the advancement of artificial intelligence, an increasing number of machine learning techniques are being applied in the field of assisted and automated driving [1][2]. Among them, Reinforcement Learning (RL) is a branch of machine learning that enables intelligent agents to learn optimal or near-optimal behaviors by interacting with their environments and receiving feedback in the form of rewards or penalties [3]. RL showcases robust learning and adaptability, particularly in real-time decision-making scenarios prevalent in automated driving [4, 5, 6].
@@
  • Perfecting RL agents to mirror human driving while ensuring safety and efficiency remains a complex task.
@@
To solve these challenges, a shift in approach is emerging with In-context Learning (ICL), [10]. In recent years, In-Context Learning (ICL) has gained substantial interest and attention within the research community. ICL represents the ability of large language models to perform tasks by interpreting examples or instructions provided in the input prompt, without requiring updates to the model’s parameters. Unlike traditional training methods, ICL does not rely on large amounts of training data or parameter updates. Instead, it uses a small amount of contextual information and examples to help the model make decisions. In the context of automated driving, ICL offers a transformative solution by enabling LLMs to interpret natural language cues and generate reward signals that guide the behavior of reinforcement learning agents. This approach hypothesizes that having been trained by a large amount of content with a wide range of topics, including driving knowledge, LLMs were endowed with the ability to model a diverse set of learned concepts and the user could use the prompt to “locate” a previously learned concept [11]. By specifying an objective with a natural language prompt, LLM can convert user goals into reward signals and intuitively guide RL agents to desired outcomes [12]. This innovative interaction between LLMs and RL agents not only advances opportunities in fields like automated driving, but also highlights the versatility of LLMs in complex decision making.
Building on the previously described theory, we present our approach to leveraging the synergy between LLMs and RL agents in automated driving scenarios. Specifically, we examine how LLM-guided reward signals influence RL training and performance, enabling the development of diverse and adaptive driving behaviors. Our proposed framework combines LLMs and RL to achieve more flexible and human-like decision-making in automated driving. By incorporating human-optimized reward designs, the RL agent learns driving behaviors that closely mimic human tendencies. This approach addresses complex decision-making challenges in dynamic, real-world scenarios, enhancing the adaptability and responsiveness of self-driving systems. The integration of LLMs and RL thus offers a promising pathway for advancing automated driving technologies. The contributions of this paper are summarized as follows:
  • We proposed an RL training framework incorporating LLMs to simplify the RL-AD agents’ reward design in AD scenarios, quoted as LLM-RL agent in this paper.
@@
The remainder of this paper is organized as follows. An overview of the related work is provided in Sections II. Section III presents the problem formulation followed by describing the proposed approach in SectionIV. The experiments and results are discussed in Sections V and VI and refsec:evaluation. Finally, the conclusion and future work are summarized in Section VII.
@@ II-A Large Language Models in Reinforcement Learning
Large Language Models (LLMs), known for their capacity to learn from minimal context and deliver human-like responses following training with human data [13], are increasingly considered as a potential solution when combined with Reinforcement Learning (RL) agents to perform higher-level tasks. In some studies, RL is employed to fine-tune existing LLMs[14, 15, 16]. Conversely, other research integrates RL with LLMs to execute specific tasks or develop particular skills[17, 18, 19]. In contrast, our research diverges from these methods by using LLMs as proxy rewards, enhancing RL model training as described in[20].
The research landscape in using LLMs as proxies for RL rewards encompasses varied approaches. Kwon et al. pioneered the use of LLMs for learning in contexts such as the Ultimatum Game and negotiation tasks, focusing solely on LLM-generated rewards. [12]. In contrast, our study employs a diverse set of rewards in dynamic automated driving system scenarios, broadening the scope of possibilities. Other researchers, like Xie et al. and Ma et al. [21][22], have explored reward design through direct Python code generation. Meanwhile, Song et al. in [23] required human inputs to iteratively fine-tune rewards, whereas another study [24] used LLM outputs as auxiliary rewards. Unlike these approaches, our method directly employs LLM responses to generate reward signals, simplifying the process and eliminating the need for human intervention.
@@ II-B Large Language Models in Automated Driving
Utilizing Large Language Models in automated driving enhances decision-making by leveraging their vast knowledge and reasoning capabilities. Research in vehical planning and control primarily focuses on two approaches: prompt engineering [25, 26] and fine-tuning pre-trained models [27, 28].
To address closed-loop driving tasks, Fu et al.  [26] utilize LLMs to comprehend and interact with environments. Wen et al.  [25] extended this framework by incorporating reflection modules to aid LLMs in decision-making. Our approach simplifies this process by considering only the ego car’s most recent action. Similarly, Cui et al.  [29] proposed a method that provides the LLM with comprehensive environmental information collected from various vehicle modules, ensuring safer results from LLM decisions. DHowever, these methods often suffer from real-time latency issues.
Other researchers have focused on fine-tuning the pre-trained model. For example, Chen et al.  [28] fine-tuned LLMs using 10,000 generated driving scenarios, while DriveMLM [30] combined linguistic processing with vehicle control to develop a more advanced system. In contrast, our work leverages LLMs to enhance RL performance, meeting real-time requirements by dynamically querying LLMs based on the current driving scenario.
@@ III-A Markov Decision Process Formalization
An automated driving scenario can be modeled as a Markov Decision Process: $M=\langle S,A,\rho,R,\gamma\rangle$, where $S$ is the state space representing various driving scenarios (e.g., traffic density, relative speeds of vehicles, lane positions, etc.). $A$ is the action space representing the set of all possible driving actions (e.g., accelerate, decelerate, change lanes, etc.). $\rho:S\times A\times S\rightarrow[0,1]$ is the transition probability function representing the likelihood of transitioning from one state to another given an action. $R:S\times A\rightarrow\mathbb{R}$ is the traditional reward function mapping states and actions to real-valued rewards. $\gamma$ is the discount factor, indicating the agent’s consideration for future rewards. The RL process can be defined within a Markov Decision Process (MDP) framework:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid\pi\right]$ (1)
where $\pi$ is the policy can be seen as the agent’s brain or the decision-making strategy of the agent or the mapping from states or observations to actions.
Inspired by [12], we employed an LLM as the proxy for reward calculations in our framework to achieve a more interpretable and adaptable reward mechanism.
@@
As illustrated in Fig. 1 our integrated framework uses the LLM to interpret textual prompts $\rho$ as input and outputs a reward signal. The text prompt $\rho$ consists of three major components: Task Description $\rho_{1}$, User Objective $\rho_{2}$ and Last Outcome $\rho_{3}$. The LLM is modeled as a function LLM: $A*\rightarrow A*$, taking the concatenated prompt $\varrho$ as input and outputting a string. A parser $A*\rightarrow{0,1}$ is then used to map the textual output of the LLM to a binary reward signal, which refers to the (2) of Fig. 1. Then, the binary reward signal will be sent to the RL agent for further training, which refers to the (3) of Fig. 1. This framework replaces the traditional reward function $R$ with a proxy reward function LLM and can be used with any RL training algorithm. In this framework, LLMs are used to create a proxy reward function:
$R_{\text{LLM}}(s,a)=f_{\text{LLM}}(\text{prompt}(s,a))$ (2)
@@ III-B Reward Function Shaping Strategies
To guide the agent’s behavior, we introduce a combined reward function that integrates various factors contributing to safety, efficiency, and LLM-based evaluation. The total reward function is expressed as:
$R_{\text{total}}(s,a)=\alpha R_{\text{safety}}(s,a)+\beta R_{\text{efficiency}}(s,a)+\gamma R_{\text{LLM}}(s,a)$ (3)
where $\alpha,\beta,$ and $\gamma$ are weights that balance the contributions of each reward component, and their sum is normalized to 1. $\alpha$ + $\beta$ + $\gamma$ = 1,  where $\alpha$, $\beta$, $\gamma\geq 0$. Each category is defined in detail as follows:
Safety Reward ($R_{\text{safety}}$)
The safety reward focuses on factors that promote safe driving behavior, such as lane discipline and collision avoidance. It is defined as:
$\begin{split}R_{\text{safety}}(s,a)=&\ \lambda_{1}\cdot\text{Lane\_discipline}(s,a)\\ &+\lambda_{2}\cdot\text{Collision\_avoidance}(s,a)\end{split}$ (4)
@@
Efficiency Reward ($R_{\text{efficiency}}$)
The efficiency reward aims to optimize fuel consumption and maintain appropriate speed. It is defined as:
$\begin{split}R_{\text{efficiency}}(s,a)=&\ \mu_{1}\cdot\text{Fuel\_efficiency}(s,a)\\ &+\mu_{2}\cdot\text{Speed\_maintenance}(s,a)\end{split}$ (5)
@@
LLM Reward ($R_{\text{LLM}}$)
The LLM reward leverages the evaluation provided by a large language model, which assesses the appropriateness of the agent’s actions based on human-like driving patterns. This reward is binary (either 0 or 1), defined as:
$R_{\text{LLM}}(s,a)=\nu\cdot\text{LLM\_feedback}(s,a)$ (6)
[…]
Adjusted Total Reward
We use a tangent function to scale the combined reward to the [0,1] range:
$\begin{split}R_{\text{ad}}(s,a)=\sigma\Big{(}&\alpha R_{\text{safety}}(s,a)\\ &+\beta R_{\text{efficiency}}(s,a)\\ &+\gamma R_{\text{LLM}}(s,a)\Big{)}\end{split}$ (7)
where $\sigma$ is the scaling function ensuring the reward remains within the [0,1] range. This adjusted reward is then used to guide the agent’s decision-making process.
Environment Reward
In addition to the main reward components, we include an environment-based reward that encourages speed optimization while penalizing collisions:
[…]
where $v$ is the current speed of the ego-vehicle, $v_{\text{min}}$ and $v_{\text{max}}$ are the minimum and maximum speed limits, respectively. $a$ and $b$ are coefficients for speed and collision penalties.
@@ IV Proposed Approach
A standard Deep Q Learning (DQN) model is set up as a baseline agent, trained with reward function 7 that are calculated based on the current driving scenario, which serves as a foundation of our proposed approach.
@@
These prompts distinctly outline the scenario as well as instruction for LLM to make decision for further steps. Scenario information included are: speed of the vehicle in control (the LLM-RL agent), speed and positions of nearby vehicles, LLM-RL agent’s last action and possible actions to choose from. Further instructions including driving rules, caution points, and specified format needed for generating and presenting responses.
@@ IV-A Basic Driving Rules and Task Description
To illustrate the context for the GPT agent, we provide a series of predefined prompts:
  • SYSTEM_MESSAGE: Specifies LLM’s role and sets the scenario’s context.
  • Driving_RULES: Provides the general safe-driving guidelines which shape the behavior pattern for the driving styles of the agents.
@@
These predefined messages and rules can serve as a foundation for the behaviour of the driving assistant. The construction and phrasing of previous descriptions are based on the work of drive-like-human [32]
@@ IV-B Context Description
To illustrate the context of the current road situation, two basic classes are defined: Vehicle and Lane. The Vehicle class encapsulates properties including speed in both directions and positional data, while the Lane class incorporates attributes including position, unique identification, and details regarding adjacent lanes.
@@ IV-D Formatting Output
This part of the prompt aims at formatting an output string that is interpretable to our program during runtime. We specify the outputs of the agent to generate only the following five actions of the ego vehicle: LANE_LEFT, IDLE, LANE_RIGHT, FASTER, and SLOWER. The LLM interprets all information above context and generates feedback. A procedure to generate a final reward is discussed as follows: We initialize Highway_Env and input a concatenated prompt, which includes the last action, the pre-define prompt, and the scene description. The last action is derived from the previous decision made by the agent. This prompt enables the agent to understand the scenario it faces and ensures that the responses generated by the agent adhere to user requirements.
@@
V Experiments
Figure 3: Meta prompt case study
@@ V-A Highway AD as a case study
The practical capabilities of leveraging the in-context learning ability of LLM in AD are illustrated through highway driving scenarios. This case study also explores the potential for AD agent personalization in highway scenarios by establishing two groups with different meta-prompts.
Fig. 3 shows an example of prompt case, in this scenario, ego car is driving in lane 1 on a highway. There is one vehicle in its current lane. The right lane change is deemed safe by the safety assessment. The previous decision the ego car made, was to stay idle, which means the ego car is in the process of following the front car. This scenario featured two distinct driving agents: the "aggressive" and "conservative" agents. A comprehensive analysis was conducted under controlled conditions to compare their performance.
Each RL agent’s driving behavior was shaped through user instructions, establishing their respective driving style within the scenario as shown in Fig. 4. Agent’s decisions show strong consistency in humanized behavior reflecting the behavior pattern meta-prompt. The agent considers several factors: Safety, Efficiency and Driving Rules. For an aggressive agent, the user has stated that efficiency is of higher priority, hence the agent is considering lane right for higher efficiency. For a conservative agent, the user has stated that safety is of higher priority, hence the agent is considering slowing down to keep distance from the leading car.
@@
Compared with manually designing a reward function, this manner is simpler and has demonstrated its effectiveness in guiding the behavior of an agent.
These observations indicate that LLMs, such as GPT in our case, offer a potential cost-efficiency alternative to performing similar roles with human feedback in certain contexts. LLMs can leverage vast amounts of text data to learn implicitly from human-generated content, interactions, and feedback, significantly reducing the human workload in reward design while achieving similar effects.